OpenStudy (anonymous):

If the period of a pendulum is 2.0 seconds and the acceleration due to gravity is 9.8 meters/secondÆ, what is the length of the cord on the pendulum?

OpenStudy (rock_mit182):



OpenStudy (rock_mit182):

(T) it is the time taken for a particle such as the one at A to move through one complete cycle, starting from point A and then back to A. the period(T) is the number of seconds per cycle

OpenStudy (rock_mit182):

\[T =2\pi \sqrt{\frac{ L }{ g }}\]

OpenStudy (rock_mit182):

solve for L and plug your values T and g